OpenStudy (anonymous):

A box with a square and open top must have a volume of 4000cm^3. Find the dimensions of the box that minimizes the amount of materials used.

HanAkoSolo (jamierox4ev3r):

remember @esam2 that the formula for volume of a rectangular prism is as follows: \(\Huge{V}=\) \(\Huge{w}\times{h}\times{l}\)

HanAkoSolo (jamierox4ev3r):

V represents volume, w represents weight, h represents height, and l represents length

OpenStudy (anonymous):

Yes, I remember . w=width

HanAkoSolo (jamierox4ev3r):

yes :) so if you keep these numbers as even as possible, then you will lessen the amount of materials needed. So we're here to figure out the dimensions to make this possible

OpenStudy (anonymous):

okay

HanAkoSolo (jamierox4ev3r):

So this is an optimization problem :) let the side of the square base=s Let the height of the box = h Then the volume of the box, V=s^2h The area of the box is s^2 There are 4 sides, and each side has an area of sh So total area of 4 sides is equivalent to 4sh If you add everything that we've found already, 4 sides + base=s^2 + 4sh. you following me so far? @esam2 ?

OpenStudy (anonymous):

I think so..

HanAkoSolo (jamierox4ev3r):

alright. Let A = s^2 + 4sh So in the problem, we're told that the volume is 4000 cm^3, which we established equals s^2h, and A = s^2 + 4sh Therefore, h= 4000 ÷ s^2 : so A= s^2 + 4(4000÷ s^2) = s^2 + 16,000 ÷ s Therefore, h= 4000 divided by s^2: so the area equals s^2 + 4( 4000/s^2) which equals s^2 + 16,000 divided by s ***i.e. A= s^2 +16,000÷ s

HanAkoSolo (jamierox4ev3r):

yikes, the glitch is real. Basically, you should be able to find that the area is equivalent to s^2 +16,000/s \[A= s ^{2} + 16,000\div s\] ^Wrote it out more clearly ^_^

HanAkoSolo (jamierox4ev3r):

using this, we can find the minimum area

OpenStudy (anonymous):

Oh okay, I get it! Thanks for your help! :)